# Dr.Mi-Bench: A Modular-integrated Benchmark for Scientific Deep Research Agent

### v1

@@ A.10 Further Related Work
Early work on LLM-based search agents established the foundation for combining reasoning with tool use. ReAct (yao2023react) demonstrated how models can interleave reasoning traces with external search actions, while subsequent systems such as Search-o1 (li2025search), Search-R1 (jin2025search), RAG-Gym (xiong2025rag), ReSearch(wang-etal-2024-searching), Collab-RAG (xu2025collab), R1-Searcher (song2025r1), and DynaSearcher (hao2025dynasearcher) expanded this paradigm with reinforcement learning, collaborative retrieval, dynamic search policies, and evaluation frameworks. These methods are effective for multi-hop question answering, but they remain limited to relatively short contexts and cannot analyze or synthesize insights from large collections of research papers or long-form articles (zhang2025web). To address more complex research-oriented scenarios, several commercial systems have recently emerged, including Gemini DR (gemini-dr), OpenAI DR (openai-blog), Perplexity DR (perplexity-dr), and Grok (grok-dr), each positioned as an end-to-end deep-research assistant. In practice, these systems follow a similar operational paradigm: they first decompose user queries into sub-goals, then invoke search engines or RAG pipelines to collect relevant evidence, and finally synthesize an answer with a textual explanation. Despite their utility, such systems remain largely opaque, and their evaluation typically reports only aggregate success without disentangling planning, retrieval, and reasoning capabilities.
@@
Several recent benchmarks have attempted to measure aspects of DR, but they remain partial in scope. GAIA (mialon2023gaiabenchmarkgeneralai) focuses primarily on short-form answer generation, providing little insight into planning or evidence attribution. BrowseComp (wei2025browsecompsimplechallengingbenchmark) and its multilingual variant BrowseComp-ZH (zhou2025browsecompzhbenchmarkingwebbrowsing), together with Humanity’s Last Exam (HLE) (phan2025humanitysexam), emphasize the retrieval stage, evaluating how well models can locate information in web documents, but they do not capture the full cycle of task decomposition, retrieval, and reasoning. More recently, the Deep Research Benchmark (futuresearch2025deepresearchbenchevaluating) has been introduced to provide a broader evaluation of research agents. While this benchmark expands the task set and evaluation dimensions compared to earlier efforts, it still operates largely at the end-to-end level and mainly tests search capability, leaving open the question of how to disentangle planning, retrieval, and reasoning performance in a transparent manner.
A.11 LLM Usage
We acknowledge the use of a large language model (LLM) to assist in the editing and polishing of this manuscript. The LLM was utilized to refine grammar, improve sentence structure, and ensure the clarity and readability of the text. All core ideas, analyses, and conclusions presented in this paper were conceived by the human authors, who retain full responsibility for the final content of this work.
A.12 Prompts
@@
A.13 Potential Paper Selection Bias
We acknowledge that our citation filter ($\geq$10 citations in 2 years) represents a deliberate trade-off, prioritizing signal-to-noise ratio over the inclusion of the “long tail”. Given the explosive growth in publications, this criterion acts as a vital heuristic to ensure the benchmark consists of high-impact, peer-validated work. Our goal was to first establish robust agent performance on research that reflects clear methodologies and community consensus, rather than risk noise from unvetted papers. We agree this is a limitation and identify generalization to the “long tail” as a key direction for future work.
A.14 Toolkit Setting
@@ A.15 Metrics Definition
Definition 6 (True Positives)
@@
Definition 7 (True Negatives)
@@
Definition 8 (False Positives)
@@
Definition 9 (False Negatives)
@@
A.16 LLM-as-a-Judge Prompts
We show the prompts used in LLM-as-a-Judge.
Prompt for LLM-as-a-Judge Evaluate the quality of the given scientific report to the question. Rate the response on three dimensions: accuracy, coverage, and redundancy. Use a scale from 1 (worst) to 10 (best). Accuracy: Assess how well the response addresses the question and provides correct information. Coverage: Evaluate whether the key points are covered in the response. Redundancy: Check for duplicate information. Please be strict in your evaluation, ensuring that scores reflect subtle differences in quality. Aim to distribute scores more evenly across the scale, avoiding clusters at the high end. Consider the following criteria for scoring: A score of 1-3 indicates major flaws in multiple dimensions. A score of 4-6 indicates moderate issues or inconsistencies. A score of 7-8 reflects generally good quality with minor flaws. A score of 9-10 is reserved for exemplary responses that excel in all dimensions. Question: {RESEARCH QUESTION} Response: {REPORT}
@@ A.16 LLM-as-a-Judge Prompts
A.17 Retrieval Matching Methodology
@@
[…]
A.18 LLM-Judge Reliability and Validation
@@
A.19 Domain Robustness and Difficulty Characterization

### v2

@@ Appendix B Further Related Work
Early work on LLM-based search agents established the foundation for combining reasoning with tool use. ReAct (Yao et al., 2023) demonstrated how models can interleave reasoning traces with external search actions, while subsequent systems such as Search-o1 (Li et al., 2025a), Search-R1 (Jin et al., 2025a), RAG-Gym (Xiong et al., 2025), ReSearch(Wang et al., 2024), Collab-RAG (Xu et al., 2025), R1-Searcher (Song et al., 2025), and DynaSearcher (Hao et al., 2025) expanded this paradigm with reinforcement learning, collaborative retrieval, dynamic search policies, and evaluation frameworks. These methods are effective for multi-hop question answering, but they remain limited to relatively short contexts and cannot analyze or synthesize insights from large collections of research papers or long-form articles (Zhang et al., 2025b). To address more complex research-oriented scenarios, several commercial systems have recently emerged, including Gemini DR (Google, ), OpenAI DR (OpenAI, 2025), Perplexity DR (PerplexityAI, ), and Grok (xAI, ), each positioned as an end-to-end deep-research assistant. In practice, these systems follow a similar operational paradigm: they first decompose user queries into sub-goals, then invoke search engines or RAG pipelines to collect relevant evidence, and finally synthesize an answer with a textual explanation. Despite their utility, such systems remain largely opaque, and their evaluation typically reports only aggregate success without disentangling planning, retrieval, and reasoning capabilities.
@@
Several recent benchmarks have attempted to measure aspects of DR, but they remain partial in scope. GAIA (Mialon et al., 2023) focuses primarily on short-form answer generation, providing little insight into planning or evidence attribution. BrowseComp (Wei et al., 2025) and its multilingual variant BrowseComp-ZH (Zhou et al., 2025), together with Humanity’s Last Exam (HLE) (Phan et al., 2025), emphasize the retrieval stage, evaluating how well models can locate information in web documents, but they do not capture the full cycle of task decomposition, retrieval, and reasoning. More recently, the Deep Research Benchmark (FutureSearch et al., 2025) has been introduced to provide a broader evaluation of research agents. While this benchmark expands the task set and evaluation dimensions compared to earlier efforts, it still operates largely at the end-to-end level and mainly tests search capability, leaving open the question of how to disentangle planning, retrieval, and reasoning performance in a transparent manner.
Appendix C Prompts
C.1 Evaluation Prompts
@@
C.2 LLM-as-a-Judge Prompts
We show the prompts used in LLM-as-a-Judge.
Prompt for LLM-as-a-Judge Evaluate the quality of the given academic report to the question. Rate the response on three dimensions: accuracy, coverage, and redundancy. Use a scale from 1 (worst) to 10 (best). Accuracy: Assess how well the response addresses the question and provides correct information. Coverage: Evaluate whether the key points are covered in the response. Redundancy: Check for duplicate information. Please be strict in your evaluation, ensuring that scores reflect subtle differences in quality. Aim to distribute scores more evenly across the scale, avoiding clusters at the high end. Consider the following criteria for scoring: A score of 1-3 indicates major flaws in multiple dimensions. A score of 4-6 indicates moderate issues or inconsistencies. A score of 7-8 reflects generally good quality with minor flaws. A score of 9-10 is reserved for exemplary responses that excel in all dimensions. Question: {RESEARCH QUESTION} Response: {REPORT}
Appendix D Potential Paper Selection Bias
Bias
Our citation filter ($\geq$10 citations in 2 years) represents a deliberate trade-off, prioritizing signal-to-noise ratio over the inclusion of the“long tail”. Given the explosive growth in publications, this criterion acts as a vital heuristic to ensure the benchmark consists of high-impact, peer-validated work. Our goal was to first establish robust agent performance on research that reflects clear methodologies and community consensus, rather than risk noise from unvetted papers. We agree this is a limitation and identify generalization to the "long tail" as a key direction for future work.
Non-STEM and Applied Scope
Beyond traditional core sciences, the benchmark explicitly incorporates non-STEM and applied domains, specifically Finance, Medicine, and Building and Construction, to ensure broad generalization capabilities.
Cross-Disciplinary Nature
We prioritized domains that inherently bridge multiple specialties (e.g., Materials and Finance) to assess the agents’ capacity for interdisciplinary knowledge synthesis.
Appendix E Toolkit Setting
@@ Appendix F Metrics Definition
Definition F.1 (True Positives).
@@
Definition F.2 (True Negatives).
@@
Definition F.3 (False Positives).
@@
Definition F.4 (False Negatives).
@@
Appendix G Retrieval Matching Methodology
@@
Appendix H Annotation Details
H.1 Annotation Statistics and Distribution
Model Accuracy Precision Recall F1
Retrieval
Post-Cutoff 25.27
Full 25.18
Reasoning
Post-Cutoff 59.41 60.18 39.34 46.66
Full 59.94 59.30 39.74 46.56
Model Accuracy Precision Recall F1
Retrieval
Post-Cutoff 37.15
Full 36.77
Reasoning
Post-Cutoff 63.97 60.97 43.33 49.91
Full 59.76 60.67 40.48 47.66
H.2 Annotation Guideline
H.3 Annotation Examples
An example of the summary writing.
An example of the research task annotation.
An example of the review task annotation.
Appendix I Do Agents Overestimate Their Capabilities through Data Leakage?
Forward-citation Leakage.
Forward citations (papers published after the source) could distort evaluation metrics by artificially lowering retrieval scores or inflating reasoning capabilities through information leakage. To address this, we implement a temporal filtering experiment by explicitly adding publication date constraints to the agent prompts. As shown in Table 9 (detailed in Table 6), the overall performance shift is minimal (Retrieval Accuracy from 25.18% to 25.27%), confirming that forward citation leakage has a negligible impact and validating the robustness of our results.
Cut-off Dates Leakage.
Appendix J Annotation Quality Management
Pre-annotation Training and Pilot Study.
To ensure high-quality and consistent annotations, all domain experts underwent a training and calibration phase. Annotators were first required to analyze a shared set of pilot papers and produce summaries based on a standardized rubric (Appendix H.3). This exercise, followed by a cross-review, served to align their understanding of what constitutes a paper’s core claims, ensuring a shared basis for all annotation tasks.
In-process Dual Annotation and Adjudication.
Our core quality control mechanism is a rigorous protocol of dual annotation and adjudication, which was applied to all four annotation stages (Query, Plan, Evidence, and Diagnostics). For every data instance, two domain experts independently produced an annotation. A third, senior annotator then compared the two versions and served as the final adjudicator, resolving any disagreements through structured discussion to produce a single consensus artifact. We computed inter-annotator agreement before adjudication and achieved a high level of consistency.
Post-annotation Validation and Maintenance.
Appendix K Further Analysis of Experiments
K.1 LLM-Judge Reliability and Validation
@@ Appendix I Do Agents Overestimate Their Capabilities through Data Leakage?
K.2 Domain Robustness and Difficulty Characterization